OpenStudy (anonymous):

Find the area of the circle in simplest polynomial form. A = 3.14r^2 A. 6.28x^2 + 3.14 B. 12.56x^2 + 3.14> C. 6.28x^2 + 3.14x + 3.14 D. 12.56x2 + 12.56x + 3.14

OpenStudy (bahrom7893):

what's r?

OpenStudy (anonymous):

This is R: 

OpenStudy (bahrom7893):

A = 3.14*(2x+1)^2 = 3.14*(4x^2 + 4x + 1) = 12.56x^2 + 12.56x + 3.14

OpenStudy (anonymous):

thnk u